Two trains are exactly 330 miles apart. They travel toward each other and meet in 3 hours. The speed of the trains differ by 20 mph. What are the speed of the trains?
Answer by algebrahouse.com(1659)

Distance = rate x time
D = rt

one train
rate = r {unknown}
time = 3 {they met in 3 hours}
d = 3r {distance = rate x time}

other train
rate = r + 20 {their speeds differ by 20}
time = 3 {they met in 3 hours}
d = 3(r + 20) = 3r + 60 {distance = rate x time, used distributive property}

3r + 60 + 3r = 330 {their combined distances, together, equals 330}
6r + 60 = 330 {combined like terms}
6r = 270 {subtracted 60 from each side}
r = 45 {divided each side by 6}
r + 20 = 65 {substituted 45, in for r, into r + 20}

the rate of one train is 45 mph
the rate of the other train is 65 mph

For more help from me, visit: www.algebrahouse.com